Grade 12General Physics

A solid cone and a hemisphere have a same base. Calculate the ratio of the height of cone and the radius of hemisphere so that the centre of mass of the common structure coincides withe centre of the base. Pls give detailed solution.

Mansi
10 Years agoGrade 12

1 Answer

Vikas TU
10 Years ago
let h be the height of cone then,
ycom of solid cone is h/4 (from base)
ycom of solid hemisphere is: 3r/8 (from base)
 
Given com for common structure should be at the centre of the base.
.i.e. (0,0)
 
Ycom =  (Volumeof cone*ycom of cone + Volumeof Hemisphere*ycom of hemisphere)/Total Volume.
 0  =  (pi*r^2*h*h/4 + 2/3*pi*r^3*3r/8)/Total Volume)
Or 
 
 (pi*r^2*h*h/4 + 2/3*pi*r^3*3r/8)  = 0
h^2/4 + 2/3*r^2*3/8 = 0
h^2/4 + r^2/4 = 0
h^2 + r^2 = 0
h^2 = -r^2
or
|h|^2 = |r|^2
(|h|/|r|)^2 = 1
or
|h|/|r| = 1
or
h/r = 1
 
hheight and radius can’t be negative.